### v1

@@ Abstract
Understanding and predicting the properties of inorganic materials is crucial for accelerating advancements in materials science and driving applications in energy, electronics, and beyond. Integrating material structure data with language-based information through multi-modal large language models (LLMs) offers great potential to support these efforts by enhancing human–AI interaction. However, a key challenge lies in integrating atomic structures at full resolution into LLMs. In this work, we introduce MatterChat, a versatile structure-aware multi-modal LLM that unifies material structural data and textual inputs into a single cohesive model. MatterChat employs a bridging module to effectively align a pretrained machine learning interatomic potential with a pretrained LLM, reducing training costs and enhancing flexibility. Our results demonstrate that MatterChat significantly improves performance in material property prediction and human-AI interaction, surpassing general-purpose LLMs such as GPT-4. We also demonstrate its usefulness in applications such as more advanced scientific reasoning and step-by-step material synthesis.
@@
Foundation model, Large Language Model, Multi-Modal Learning, Inorganic Material1 Introduction
@@
This bottleneck has inspired a wave of exploration into how natural language processing (NLP), and particularly large language models (LLMs), might be leveraged to fill the gap. LLMs like BERT [15], GPT [16], and newer open-source LLMs such as Mistral [17], Llama [18] and DeepSeek [19] have shown substantial promise across different domains by interpreting and generating language. Trained on extensive datasets, these models can support some scientific tasks that require interpretive language capabilities, such as question-answering (QA) [20] and retrieving information from unstructured text sources [21]. In recent years, there has been several efforts incorporating LLM to solve material related problems [22, 23], either by leveraging pretrained LLMs or multi-modal LLMs. Unfortunately, these approaches rely on text-based representations, such as chemical formula [22], Simplified Molecular Input Line Entry System (SMILES) strings [24, 23], textual descriptions [25], or Crystallographic Information File (CIF) [26], which lose the full resolution of atomic structures. Consequently, they exhibit inferior performance compared to pure graph-based models for predicting material properties [25], and they also potentially hinder other downstream tasks where structural information is crucial. Thanks to the steady development of MLIPs [5, 6, 7], particularly in their universal form (uMLIPs) [7], these models can now serve as atomistic pre-trained models capable of supporting a wide range of applications. The locality assumption underlying uMLIPs ensures that they effectively represent the local environment of each atom. Therefore, it is feasible to extract structural information from atom embeddings in a pretrained uMLIP.
@@ 2.1 Overview of MatterChat
Material Structure Branch. The Material Structure Branch encodes material structures as graphs that capture the atomic local environment. To process these graphs, we use CHGNet [27], a state-of-the-art graph-based uMLIP model designed for crystal structures, to process these graphs. CHGNet is pretrained on a diverse dataset of materials, encompassing a wide range of symmetries, compositions, and bonding types, enabling it to effectively model complex atomic interactions and structural details. By capturing essential compositional features, such as atomic types and chemical bonds, along with spatial features like bond angles, CHGNet generates high-quality atom embeddings that are both physically meaningful and well-suited for downstream tasks.
@@
Language Branch. The language branch is used to process user’s text-based prompts, such as requests for property predictions, chemical formulas, space group information, or other material characteristics. We use the Mistral 7B LLM [17], one of the latest open-source LLMs, chosen for its exceptional performance across a wide range of scientific and non-scientific tasks. This branch processes each prompt, transforming it into dense embeddings that capture the semantic content of the inquiry. These embeddings are then combined with the query embeddings processed by the bridge model using a structured fusion approach, allowing the model to effectively incorporate both textual and material information. This integration enables the LLM to generate precise and contextually relevant responses tailored to the user’s specific material-related prompts.
@@
Bridge Model. To facilitate the integration between atom embeddings and the language branch, we developed a bridge model inspired by the BLIP2 architecture [28], based on a multi-layer transformer framework. This bridge model includes 32 trainable query vectors that interact with atom embeddings using an alternating attention mechanism. Cross-attention in even-numbered layers extracts key features from the atom embeddings, while self-attention in odd-numbered layers enhances representational depth. This approach refines the atom embeddings into query embeddings that are most connected to text, as shown in Figure 1a. Finally, these refined representations are mapped to LLM-compatible embeddings via a linear projection layer.
@@
Figure 1(b)-(c) provide an overview of the dataset of crystalline structures used in our training set. Figure 1(b) visualizes the material distribution on the periodic table, highlighting that the dataset evenly spans a diverse range of elements up to Plutonium. Figure 1(c) depicts the distribution of crystalline structures by space group across the dataset. The dataset was curated from the Materials Project [29] and contains 142,899 material structures. For each structure, we generated a corresponding text-based dataset encompassing 12 tasks: 3 descriptive tasks (chemical formula, space group, and crystal system) and 9 property prediction tasks. These property prediction tasks include metallicity, direct bandgap, stability, experimental observation (Exp Ob), magnetic status (Is Magnetic), magnetic order, formation energy, energy above the hull, and bandgap, as summarized in Figure 1(a). Further details regarding the training scheme, hyper-parameters, and dataset curation are provided in the Methods section.
@@ 2.2 Multi-modal Material task interactions with MatterChat
Figure 2 illustrates the human-AI interaction examples with MatterChat, across a diverse range of material property prediction and analysis tasks. The figure shows MatterChat’s ability to effectively address a broad spectrum of user prompts ranging from fundamental material attributes (e.g., chemical formulas, space groups and crystal system) to complex material properties (e.g., thermal stability, bandgaps, formation energies and energy above the hull). The top panels in Figure 2 shows three interactive examples of material property prompts from randomly selected materials from the Material Project Dataset. The top left panel presents a human-AI query interface with MatterChat for the material identified by ID “mp-1001021”. It provides a detailed profile including the chemical formula , its crystalline structure denoted by the space group Fd-3m, and electronic properties such as a bandgap of 0.21350 eV. The interface also addresses the material’s lack of thermal stability. The top mid panel shows the interaction example with the material with mp ID of “mp-1028281”. It provides a comprehensive breakdown of the material’s composition attributes, including its chemical formula (Mg14V4Sb) and its space group (Amm2). The interaction further predicts that the material is both magnetic and metallic, and its formation energy is estimated at 0.05912 eV/atom. The top right panel provides an interaction example with MatterChat of the material with Materials Project ID of “mp-10198”. This panel informs user’s query about the chemical composition Mn3PdN and its cubic crystal structure, with the space group classified as Pm-3m. Additionally, it estimated that the material possesses an indirect bandgap, which is an important characteristic for applications in electronics. MatterChat also accurately predicts the ferromagnetic (FM) magnetic behaviors that the material exhibits, and it mentions its energy above hull value at 0.03571 eV. In the bottom panel, we present a comparative evaluation of MatterChat’s performance on formation energy evaluation tasks for newly discovered materials from GNoME [30]. The model was compared against commercial LLMs, Gemini [31], GPT-4o [32] and DeepSeek [19]. The results show MatterChat’s superior accuracy in estimating formation energies, consistently delivering predictions closer to the ground truths. For example, MatterChat’s formation energy predictions for “mp-3202380” and “mp-3206774” show a remarkable alignment with the ground truth values. These results demonstrate MatterChat’s ability to integrate structural and textual data seamlessly for a wide range of material property tasks.
@@
Figure 3 demonstrates MatterChat’s advanced reasoning capabilities, showing how it leverages the comprehensive knowledge base of LLMs to address complex materials science challenges. By using a multi-modal query system, MatterChat effectively combines material structure data with textual reasoning. This integration facilitates a working memory scheme [33], which enables the model to provide domain-specific reasoning, detailed synthesis procedures, and explanations that are deeply grounded in the structural properties of materials. Panel (a) presents the chat log for silicon with the space group of i43m. MatterChat not only retrieves the chemical formula and correct space group, but it also provides a rationale for the structural instability of this silicon phase. The model explains that the i43m space group exhibits higher energy per unit cell compared to the thermodynamically stable cubic diamond structure of silicon, making it less likely to occur under standard conditions. Panel (b) illustrates an interaction regarding a popular semiconductor material Gallium Nitride (GaN). Here, MatterChat accurately identifies the chemical formula and space group (P63mc) and constructs a detailed synthesis procedure that aligns with established methods such as molecular beam epitaxy (MBE), metal-organic chemical vapor deposition (MOCVD), and hydride vapor phase epitaxy (HVPE), demonstrating the model’s ability to apply structural and contextual knowledge in generating practical scientific outputs. Figure 3(c) explores an interaction for a widely used ferrite material, Yttrium Iron Garnet (YIG, ). MatterChat is able to take the structure and generate detailed text description. Additionally, MatterChat can further generate the synthesis protocol from the preparation of initial materials, their precise mixing ratios, the required sintering conditions, and subsequent characterization techniques, such as X-ray diffraction (XRD) and scanning electron microscopy (SEM). This example demonstrates MatterChat’s capability to synthesize and apply domain-specific knowledge effectively, aligning closely with established scientific practice [34].
@@ 2.3 MatterChat extracted embeddings contain structural and property information
We further explore MatterChat’s ability to leverage material structural information by providing a detailed visualization/clustering analysis with the UAMP dimension reduction technique [35]. Figure 4(a) to (e) show comprehensive visualizations of embeddings processed by the bridge model, with all material samples that contain silicon (Si), carbon (C), and their composites compounds (e.g., , ) from the Material Project database [36]. UMAP was used to reduce the embeddings from an original 4096 dimensions to two dimensions, with the x and y axes corresponding to the first and second reduced dimensions, respectively.
@@
Figure 4(a) presents the visualizations containing all the selected materials; each sample is color-coded with a structure similarity score [37]. The clustering generally follows distinctions in chemical compositions. Additionally, materials with the same atomic composition are grouped into separate clusters based on crystalline structural differences (e.g., Carbon with Diamond vs Graphite crystalline structure). Figures 4(b) and 4(d) show the filtered visualization/clustering result for materials have single composition of Si and SiC. Figure 4(b) shows the gradient of structure similarity scores, ranging from blue (low similarity) to red (high similarity), demonstrating how closely related structural features result in spatial proximity within the embedding space. However, an interesting exception is observed with SiC: despite its identical composition and similar structural phases, two distinct clusters of SiC emerge, suggesting that factors beyond composition and structure alone influence their separation. To further explore factors that influence clustering, we labeled the samples according to their formation energy, with results displayed in Figures 4(c) for SiC and 4(e) for Si. These figures clearly show a trend from low to high formation energy. This analysis reveals that clusters grouped by structural similarity also align closely in terms of formation energy. Such findings indicate that the model’s ability to produce embeddings that not only differentiate structural characteristics but also correlate with key material properties. Given that the embeddings derived from bridge model preserve both material structure and property-relevant information, we implemented a multi-modal Retrieval-Augmented Generation (RAG) mechanism during inference, as illustrated in Figure 4(f). Instead of relying solely on a single output from MatterChat for each query-sample pair, we now retrieve additional information of two more samples from the material pool (training set). This retrieval is based on the L2 similarity between the embeddings of the sample material and those in the pool. After that we aggregate all three result to get the final output by applying a majority-voting strategy for classification tasks and averaging for quantitative tasks. Such method could further enhance the overall robustness of MatterChat across different tasks. The details of visualization method are provided in Method Section.
@@ 2.4 Comprehensive quantitative analysis for all material tasks
To provide a more thorough evaluation of MatterChat, we conducted a comprehensive analysis of its performance across nine material property tasks on the evaluation set (14290 samples). This analysis compares MatterChat to two open-source large language models: (LLMs)—Vicuna [38], and Mistral [17], two physical ML models: SchNet [39] and CHGNet [27].
@@
Six of the nine tasks are classification tasks, evaluated using prediction accuracy as the metric. These tasks include metallicity, bandgap type, stability, magnetic properties, and experimental observables. As shown in the top two rows of Figure 5, MatterChat demonstrates improvements over all tested models. Compared to state-of-the-art LLMs, MatterChat achieves substantially higher accuracy, benefiting from its ability to integrate graph-based material structure data with natural language reasoning. Moreover, MatterChat even outperforms physical ML models such as SchNet and CHGNet, showing its superior capability in handling multi-modal inputs for material classification.
@@
The remaining three tasks—formation energy, energy above the hull, and bandgap prediction—are numerical property prediction tasks. For these tasks, the mean squared error (MSE) between predicted values and ground truth is used as the evaluation metric. Due to the inherent limitations of LLMs in producing accurate numerical predictions [40], their performance is not included in the bar plots, as the errors were excessively large. Instead, Figure 5 compares MatterChat with the physical ML models.

### v2

@@ Abstract
Understanding and predicting the properties of inorganic materials is crucial for accelerating advancements in materials science and driving applications in energy, electronics, and beyond. Integrating material structure data with language-based information through multi-modal large language models (LLMs) offers great potential to support these efforts by enhancing human–AI interaction. However, a key challenge lies in integrating atomic structures at full resolution into LLMs. In this work, we introduce MatterChat, a versatile structure-aware multi-modal LLM that unifies material structural data and textual inputs into a single cohesive model. MatterChat employs a bridging module to effectively align a pretrained universal machine learning interatomic potential with a pretrained LLM, reducing training costs and enhancing flexibility. Our results demonstrate that MatterChat significantly improves performance in material property prediction and human-AI interaction, surpassing general-purpose LLMs such as GPT-4. We also demonstrate its usefulness in applications such as more advanced scientific reasoning and step-by-step material synthesis.
@@
Foundational model, Large Language Model, Multi-Modal Learning, Inorganic Material1 Introduction
@@
This bottleneck has inspired a wave of exploration into how natural language processing (NLP), and particularly large language models (LLMs), might be leveraged to fill the gap. LLMs like BERT [15], GPT [16], and newer open-source LLMs such as Mistral [17], Llama [18] and DeepSeek [19] have shown substantial promise across different domains by interpreting and generating language. Trained on extensive datasets, these models can support some scientific tasks that require interpretive language capabilities, such as question-answering (QA) [20] and retrieving information from unstructured text sources [21]. In recent years, there has been several efforts incorporating LLM to solve material related problems [22, 23], either by leveraging pretrained LLMs or multi-modal LLMs. Unfortunately, these approaches rely on text-based representations, such as chemical formulas [22], Simplified Molecular Input Line Entry System (SMILES) strings [24, 23], textual descriptions [25], or Crystallographic Information File (CIF) [26], which lose the full resolution of atomic structures. Consequently, they exhibit inferior performance compared to pure graph-based models for predicting material properties [25], and they also potentially hinder other downstream tasks where structural information is crucial. Thanks to the steady development of MLIPs [5, 6, 7], particularly in their universal form (uMLIPs) [7], these models can now serve as atomistic pre-trained models capable of supporting a wide range of applications. The locality assumption underlying uMLIPs ensures that they effectively represent the local environment of each atom. Therefore, it is feasible to extract structural information from atom embeddings in a pretrained uMLIP.
@@ 2.1 Overview of MatterChat
Material Processing Branch. The Material Processing Branch encodes material structures as graphs that capture the atomic local environment. We use CHGNet [27], a state-of-the-art graph-based uMLIP model designed for crystal structures, to process these graphs. CHGNet is pretrained on a diverse dataset of materials, encompassing a wide range of symmetries, compositions, and bonding types, enabling it to effectively model complex atomic interactions and structural details. By capturing essential compositional features, such as atomic types and chemical bonds, along with spatial features like bond angles, CHGNet generates high-quality atom embeddings that are both physically meaningful and well-suited for downstream tasks.
@@
Language Processing Branch. The Language Processing Branch is used to process user’s text-based prompts, such as requests for property predictions, chemical formulas, space group information, or other material characteristics. We use the Mistral 7B LLM [17], one of the latest open-source LLMs, chosen for its exceptional performance across a wide range of scientific and non-scientific tasks. This branch processes each prompt, transforming it into dense embeddings that capture the semantic content of the inquiry. These embeddings are then combined with the query embeddings processed by the bridge model using a structured fusion approach, allowing the model to effectively incorporate both textual and material information. This integration enables the LLM to generate precise and contextually relevant responses tailored to the user’s specific material-related prompts.
@@
Bridge Model. To facilitate the integration between atom embeddings and the Language Processing Branch, we developed a bridge model inspired by the BLIP2 architecture [28] based on a multi-layer transformer framework. This bridge model includes 32 trainable query vectors that interact with atom embeddings using an alternating attention mechanism. Cross-attention in even-numbered layers extracts key features from the atom embeddings, while self-attention in odd-numbered layers enhances representational depth. This approach refines the atom embeddings into query embeddings that are most connected to text, as shown in Figure 1a. Finally, these refined representations are mapped to LLM-compatible embeddings via a linear projection layer.
@@
Figures 1(b)-(c) provide an overview of the dataset of crystalline structures used in our training set. Figure 1(b) visualizes the material distribution on the periodic table, highlighting that the dataset evenly spans a diverse range of elements up to Plutonium. Figure 1(c) depicts the distribution of crystalline structures by space group across the dataset. The dataset was curated from the Materials Project [29] and contains 142,899 material structures. For each structure, we generated a corresponding text-based dataset encompassing 12 tasks: 3 descriptive tasks (chemical formula, space group, and crystal system) and 9 property prediction tasks. These property prediction tasks include metallicity, direct bandgap, stability, experimental observation (Exp Ob), magnetic status (Is Magnetic), magnetic order, formation energy, energy above the hull, and bandgap, as summarized in Figure 1(a). Further details regarding the training scheme, hyper-parameters, and dataset curation are provided in the Methods section.
@@ 2.2 Multi-modal material task interactions with MatterChat
Figure 2 illustrates examples of human-AI interaction with MatterChat across a diverse range of material property prediction and analysis tasks. It shows MatterChat’s ability to effectively address a broad spectrum of user prompts ranging from fundamental material attributes (e.g., chemical formulas, space groups and crystal system) to complex material properties (e.g., thermal stability, bandgaps, formation energies and energy above the hull). Figure 2(a) shows three interactive examples of material property prompts from randomly selected materials from the Material Project Dataset. The top left panel presents a human-AI query interface with MatterChat for the material with mp-id of “mp-1001021”. It provides a detailed profile including the chemical formula , its crystalline structure denoted by the space group Fd-3m, and electronic properties such as a bandgap of 0.21350 eV. The interface also addresses the material’s lack of thermal stability. The top mid panel shows the interaction example with the material with mp-id of “mp-1028281”. It provides a comprehensive breakdown of the material’s composition attributes, including its chemical formula () and its space group (Amm2). The interaction further predicts that the material is both magnetic and metallic, and its formation energy is estimated at 0.05912 eV/atom. The top right panel provides an interaction example with MatterChat of the material with mp-id of “mp-10198”. This panel informs the user’s query about the chemical composition and its cubic crystal structure, with the space group classified as Pm-3m. Additionally, it estimated that the material possesses an indirect bandgap, which is an important characteristic for applications in electronics. MatterChat also accurately predicts the ferromagnetic (FM) magnetic behaviors that the material exhibits, and it mentions its energy above hull value at 0.03571 eV/atom. In the bottom panel, we present a comparative evaluation of MatterChat’s performance on formation energy evaluation tasks for newly discovered materials from GNoME [30]. The model was compared against commercial LLMs, Gemini [31], GPT-4o [32] and DeepSeek [19]. The results show MatterChat’s superior accuracy in estimating formation energies, consistently delivering predictions closer to the ground truths. For example, MatterChat’s formation energy predictions for “mp-3202380” and “mp-3206774” show a remarkable alignment with the ground truth values. These results demonstrate MatterChat’s ability to integrate structural and textual data seamlessly for a wide range of material property tasks.
@@
Figure 3 demonstrates MatterChat’s advanced reasoning capabilities, showing how it leverages the comprehensive knowledge base of LLMs to address complex materials science challenges. By using a multi-modal query system, MatterChat effectively combines material structure data with textual reasoning. This integration facilitates a working memory scheme [33], which enables the model to provide domain-specific reasoning, detailed synthesis procedures, and explanations that are deeply grounded in the structural properties of materials. Panel (a) presents the chat log for silicon with the space group of cmcm. MatterChat not only retrieves the chemical formula and correct space group, but it also provides a rationale for the structural instability of this silicon phase. The model explains that the cmcm space group exhibits higher energy per unit cell compared to the thermodynamically stable cubic diamond structure of silicon, making it less likely to occur under standard conditions. Panel (b) illustrates an interaction regarding a popular semiconductor material Gallium Nitride (GaN). Here, MatterChat accurately identifies the chemical formula and space group (P63mc) and constructs a detailed synthesis procedure that aligns with established methods such as molecular beam epitaxy (MBE), metal-organic chemical vapor deposition (MOCVD), and hydride vapor phase epitaxy (HVPE), demonstrating the model’s ability to apply structural and contextual knowledge in generating practical scientific outputs. Figure 3(c) explores an interaction for a widely used ferrite material, Yttrium Iron Garnet (YIG, ). MatterChat is able to take the structure and generate detailed text descriptions. Additionally, MatterChat can further generate the synthesis protocol from the preparation of initial materials, their precise mixing ratios, the required sintering conditions, and subsequent characterization techniques, such as X-ray diffraction (XRD) and scanning electron microscopy (SEM). This example demonstrates MatterChat’s capability to synthesize and apply domain-specific knowledge effectively, aligning closely with established scientific practice [34].
@@ 2.3 MatterChat extracted embeddings contain structural and property information
We further explore MatterChat’s ability to leverage material structural information by providing a detailed visualization/clustering analysis with the UAMP dimension reduction technique [35]. Figures 4(a)-(e) show comprehensive visualizations of embeddings processed by the bridge model, with all material samples that contain silicon (Si), carbon (C), and their composites compounds (e.g., , ) from the Material Project database [36]. UMAP was used to reduce the embeddings from an original 4096 dimensions to two dimensions, with the x and y axes corresponding to the first and second reduced dimensions, respectively.
@@
Figure 4(a) presents the visualizations containing all the selected materials; each sample is color-coded with a structure similarity score [37]. The clustering generally follows distinctions in chemical compositions. Additionally, materials with the same atomic composition are grouped into separate clusters based on crystalline structural differences (e.g., Carbon with Diamond vs. Graphite crystalline structure). Figures 4(b) and 4(d) show zoomed-in visualizations of clustering results for materials consisting exclusively of Si and SiC compositions. Figure 4(d) shows the gradient of structure similarity scores, ranging from blue (low similarity) to red (high similarity), demonstrating how closely related structural features result in spatial proximity within the embedding space. However, an interesting exception is observed with SiC (see Figure 4(b)): despite its identical composition and similar structural phases, two distinct clusters of SiC emerge, suggesting that factors beyond composition and structure alone influence their separation. To further explore factors that influence clustering, we labeled the samples according to their formation energy, with results displayed in Figures 4(c) for SiC and 4(e) for Si. These figures clearly show a trend from low to high formation energy. This analysis reveals that clusters grouped by structural similarity also align closely in terms of formation energy. Such findings indicate that the model’s ability to produce embeddings that not only differentiate structural characteristics but also correlate with key material properties. Given that the embeddings derived from bridge model preserve both material structure and property-relevant information, we implemented a multi-modal Retrieval-Augmented Generation (RAG) mechanism during inference, as illustrated in Figure 4(f). Instead of relying solely on a single output from MatterChat for each query-sample pair, we now retrieve additional information of two more samples from the material pool (training set). This retrieval is based on the L2 similarity between the embeddings of the sample material and those in the pool. After that, we aggregate all three results to get the final output by applying a majority-voting strategy for classification tasks and averaging for quantitative tasks. Such a method could further enhance the overall robustness of MatterChat across different tasks. The details of the visualization method are provided in the Method Section.
@@ 2.4 Comprehensive quantitative analysis for all material tasks
To provide a more thorough evaluation of MatterChat, we conducted a comprehensive analysis of its performance across nine material property tasks on the evaluation set (14290 samples). This analysis compares MatterChat with two open-source large language models (LLMs), Vicuna [38] and Mistral [17], as well as two physical ML models, SchNet [39] and CHGNet [27].
@@
Six out of the nine material property tasks are classification tasks, evaluated using prediction accuracy as the metric. These tasks include metallicity, bandgap type, stability, magnetic properties, and experimental observables. As shown in the top two rows of Figure 5, MatterChat demonstrates improvements over all tested models. Compared to state-of-the-art LLMs, MatterChat achieves substantially higher accuracy, benefiting from its ability to integrate graph-based material structure data with natural language reasoning. Moreover, MatterChat even outperforms physical ML models such as SchNet and CHGNet, the latter being the built-in model used in the Material Processing Branch, showing its superior capability in handling multi-modal inputs for material classification.
@@
The remaining three tasks—formation energy, energy above the hull, and bandgap prediction—are numerical property prediction tasks. For these tasks, the root mean squared error (RMSE) between predicted values and ground truth is used as the evaluation metric. Due to the inherent limitations of LLMs in producing accurate numerical predictions [40], their performance is not included in the bar plots, as the errors were excessively large. Instead, Figure 5 compares MatterChat with the physical ML models.